In 4 years time, a father will be 3 times as old as his son. 6 years ago he was 7 times as old as his son. How old are they now?
Answer by scott8148(6628)

"In 4 years time, a father will be 3 times as old as his son" ___ f + 4 = 3(s + 4)

"6 years ago he was 7 times as old as his son" ___ f - 6 = 7(s - 6) ___ f - 6 = 7s - 42 ___ f = 7s - 36

substituting ___ (7s - 36) + 4 = 3s + 12

4s = 12 - 4 + 36 = 44 ___ s = 11

substituting ___ f = 7(11) - 36 ___ f = 41